OpenStudy (anonymous):

find a radical equation of the form sqrt(ax+b)=x+c so that one solution is extraneous.. please help!

OpenStudy (anonymous):

there is also an addition to the question... a) Is there a value for H that makes it possible for the equation sqrt(x+h)+5=0 to have any real number solutions? Explain.

OpenStudy (anonymous):

b) Explain the relationship between the solutions to the equation sqrt(x-3)-2=0 and the graph of the function y=sqrt(x-3)-2